OpenStudy (anonymous):

explain these questions calc 2

OpenStudy (anonymous):

\[\sum_{k=1}^{\inf}(\pi/e)^{-k}\]

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

does that function converge why or why not?

OpenStudy (anonymous):

it is a geometric series, a rather silly looking one, but a geometric one in any case

OpenStudy (anonymous):

\[a+ar+ar^2+ar^3+...=\frac{a}{1-r}\] this converges because \[\frac{\pi}{e}<1\]

OpenStudy (anonymous):

not to be too pedantic, but is not a "function" it is a number

OpenStudy (anonymous):

ya, i agree but it's past midnight here.

OpenStudy (anonymous):

onto question 2 if you've got time.

OpenStudy (anonymous):

LOL ... technically, not to be pedantic, it is an infinite series, whose sum is a number!

OpenStudy (anonymous):

if L is a real number and \[\sum_{k=12}^{\inf}A ^{k} converges then \sum_{k=1}^{\inf}A ^{k}\]

OpenStudy (anonymous):

if A is a real number*